# DHP: Discrete Hierarchical Planning for HRL Agents

### v2

@@ 2.4.1 Exploratory Reward
Since the planning policy $\pi_{p}$ uses the GCSR decoder for subgoal prediction, it is limited to the subgoals learned by the GCSR module. The GCSR training data must contain all possible state triplets in the environment for the planner to function well. To do this, we formulate our exploratory objective to encourage the Explorer to enact trajectories not modeled well by the GCSR module. If Explorer traverses a trajectory that contains a state triplet $(s_{t},s_{t+q/2},s_{t+q})$, the modeling error is measured as the mean squared error between sub-goal $s_{t+q/2}$ and its prediction using the GCSR decoder. We compute the exploratory rewards for state $s_{t}$ based on the previous states as:
@@ 2.4.2 Memory Augmented Explorer
Since the exploratory rewards for the current step depend on the past states, the Explorer needs to know them to guide the agent accurately along rewarding trajectories. Fig. 5(c) shows the state dependencies for the exploratory rewards at different temporal resolutions $q$. Each temporal resolution $q$ requires $(s_{t-q},s_{t-q/2})$ for computing reward. For our case, the past states required are: $[(s_{t-K},s_{t-2K}),(s_{t-2K},s_{t-4K}),...]$ for each $q\in\{2K,4K,...\}$ respectively. Removing the duplicates reduces the set of states required to: $\{s_{t},s_{t-K},s_{t-2K},...\}$.
@@
Figure 5: (a) Unconditional VAE that learns to predict states not conditioned on other states. (b) The explorer using the memory and the VAE decoder to predict subgoals for the worker. (c) Illustration showing the states required as inputs for the Explorer for an example trajectory (shown at the top) being played out by an agent. It is a coarse trajectory that shows every $K$-th frame. The agent is at state $s_{t}$ and will receive rewards when it moves into the placeholder future state $s_{t+1}$ (dashed border). The rewards at $s_{t+1}$ will be computed using the GCSR for different temporal resolutions $q\in Q$ indicated on the right. The colored arrows indicate the state triplet required to compute the exploratory reward at $s_{t+1}$. Combining these state dependencies and removing redundancies yields the input requirements indicated below the dashed line. The inputs consist of the current state and the memory.
@@ 2.4.3 Policy Optimization
The Explorer is optimized as an SAC using the policy gradients from the REINFORCE objective. RSSM is used to imagine trajectories $\kappa$, followed by extracting every $K$-th frame of the rollout as abstract trajectories. For clarity, let the time steps of the abstract trajectory be indexed by $k$, where $k=t/K$. Then, the lambda returns $G^{\lambda}_{k}$ are computed for the abstract trajectories using a discount factor $\gamma_{L}$ (Eq. 12). Finally, the lambda returns are used to formulate the explorer actor and critic loss (Eq. 13,14).
@@ 3 Evaluation & Results
We extensively test our agent in the 25-room environment, a 2D maze task where the agent has to navigate through connected rooms to reach the target position. Benchmarks from previous methods show average episode duration $>150$ steps, indicating a long-horizon task. Observations are provided as initial and goal states ($64\times 64$ images) and a reward $0<R\leq 1$ upon reaching the goal position. Each episode lasts $400$ steps before terminating with a 00 reward. We use the same evaluation parameters as the previous benchmarks and average across $100$ runs [27]. We also test our method in momentum-based environments like RoboYoga [20], which can be challenging due to the lack of momentum information in a single image (goal).
@@
We use a common hyperparameter setup for all tasks. The goal refresh rate is set to $K=8$, and the modeled temporal resolutions as $Q=\{2K,4K,8K,16K,32K\}$. The depth of the unrolled tree during training is $D=5$ and during inference is $D_{I}=8$ unless specified otherwise. For the first $3$M steps, the explorer is used as the manager, then it shifts to the planning policy. The agent is trained every $16$ environmental steps. Please refer to section B for complete training details.
@@ 3.1 Results
Figure 7: Samples plan by our agent during training and inference. (a) Samples of full plans as sequences of subgoals from the start states (left) to the goal states (right). The subgoals are extracted from the goal states of all terminal nodes. The blanks are when nodes terminate before maximum depth. (b) Sample subgoals generated during inference. The first and last images indicate the initial and goal states. Other images represent subgoals that break the path from the initial to the image on its right [Same order as 2(b)]. Blanks are when a reachable subgoal is found before the max depth.
We compare the performance in terms of the average success rate in reaching the goal state and the average path length against previous methods. Goal-Conditioned Behavioral Cloning (GC BC, [24]) that learns goal-reaching behavior from example goal-reaching behavior. Visual foresight (VF, [9]) that optimizes rollouts from a forward prediction model via the cross-entropic method (CEM, [29, 23]). Hierarchical planning using Goal-Conditioned Predictors (GCP,[27]) optimized using CEM to minimize the predicted distance cost. Goal-Conditioned Director (GC-Director). And LEXA, a SOTA sequential planning method that also uses an explorer and a planner but optimizes continuous rewards cosine and temporal. GC BC, VF, and GCP performances are taken from [27], which uses the same evaluation strategy. Table 1 shows that our model outperforms the previous work in terms of success rate and average episode lengths. Our method and LEXA (Temp) yield the shortest episode lengths. However, note that while LEXA explicitly optimizes for path lengths, our method uses an implicit objective.
Agent Success rate Average episode length Sample complexity Time complexity
GC BC $7\%$ $402.48$ $1$ $1$
GC-Director $9\%$ $378.89\pm 87.67$ $1$ $1$
LEXA (Cos) $20\%$ $321.04\pm 153.29$ $1$ $1$
VF $26\%$ $362.82$ $MN$ $N$
@@
Table 1: Average Performance of different approaches on the 25-room navigation task over $100$ evaluation runs. $N$ is the number of plan steps, and $M$ is the number of samples generated per plan.
Fig. 8(a) shows the score and episode length plots for some methods. Here, we plot an extra experiment, Director (Fixed Goal), where the goal remains fixed and the agent only inputs the current state image. It can be seen that the agent shows signs of learning, the score falls but rises again around $8$M steps to $\sim 70\%$. Comparing this to GC-Director (which completely fails) shows that the issue is not navigation or agent size, but the complexity of a goal-conditioned long-horizon task that requires planning.
@@
Figure 8: Experimental results (mean and stddev across $3$ seeds) shown with episodic reward (top row) and average episode length (bottom row) against the environmental steps. The sharp performance rise for our methods indicates the switch from exploration to planning. (a) Baselines: Comparing our method DHP with state-of-the-art methods. (b) Bootstrapping: Compares shallow training depth models $(D=1,3)$ with the default training depth. (c) Explore data Comparison of planning policy performance trained on different data. (d) Comparison of Other reward schemes.
@@ 3.2 Ablations
We train two DHP agents with maximum tree depth during planning to $3$ and $1$. The $1$ depth agent is only allowed to break the given task once for learning. Fig. 8(b) shows the comparison of the agents, and it can be seen that all agents perform similarly.
@@
We compare the performance of the planning policy trained against three variants: using expert data (collected using a suboptimal policy in [27]), vanilla exploration objective using the reconstruction errors from the unconditional VAE (Sec. 2.4), and using an explorer without the memory. Fig. 8(c) shows the comparison, and it can be seen that the default agent performs noticeably better.
@@
Figure 9: Robo Yoga Environments
We test our agent in the Deepmind Control Suite [34] based RoboYoga [20] environment. The task requires walker and quadruped agents to reach a goal body orientation specified by an image. The environment is challenging because it rewards the agent to maintain the goal position, while our agent only plans to reach the goal state. Also, the goal states do not encode the momentum information. Fig. 9 shows the performance, of our agent at the tasks. Our agent can solve the walker task, but struggles at quadruped. We observed that our walker agent maintains the overall pose but constantly sways about the goal position. Fig. 13 shows an episode where the agent headstands constantly.
@@
We compare the default agent against a few variants: DHP (Neg Rew) That rewards $-1$ at non-terminal nodes and 00 at terminal (like an existence penalty), DHP (GAE) that estimates advantages as $\min$ of the GAE advantages for each child node, and DHP (Dist Sum) rewards negative estimated distances (between $n_{i,0}$ and $n_{i,1}$) at each node. Fig. 8(d) shows that the Neg Rew agent performs as well as the default method, while the others don’t perform as well.
@@ 4 Related Work
Hierarchical RL Agents: Hierarchical reinforcement learning (HRL) is a set of techniques that abstracts actions [4, 35, 3, 32, 26, 21]. Foundational works like the options framework [32] and MAXQ decomposition [8] introduced temporal abstraction, allowing agents to reason at multiple time scales. Modern approaches learn hierarchical policies through mutual information (Causal InfoGAN [17], DADS [30]), latent space planning (Director [14]), or trajectory encoding (OPAL [1]). These show that hierarchical decomposition aids efficient exploration and credit assignment for planning.
Planning Algorithms: Planning methods aim to solve long-horizon tasks efficiently by exploring future states and selecting optimal actions [18, 6]. Monte Carlo Tree Search (MCTS) [5] expands a tree of possible future states by sampling actions and simulating outcomes. While effective in discrete action spaces, MCTS struggles with scalability in high-dimensional or continuous environments. Visual Foresight methods [9, 11, 13] learned visual dynamics models to simulate future states, enabling planning in pixel-based environments. However, they require accurate world models and can be computationally expensive. Some use explicit graph search over the replay buffer data [10]. Model Predictive Control (MPC) [22, 23] is an online planner that samples future trajectories and optimizes actions over a finite horizon. These methods rely on sampling the future state and thus do not scale well with the horizon length. LEXA [16] is a policy-based linear planner that also uses an explorer for data collection.
To address the challenges of long-horizon tasks, some planning algorithms decompose complex problems into manageable subtasks by predicting intermediate subgoals [25, 15, 27]. MAXQ [8] decomposes the value function of the target Markov Decision Process (MDP) into an additive combination of smaller MDPs. Sub-Goal Trees [15] and CO-PILOT [2] learn a subgoal prediction policy optimized to minimize the total predicted distance measures of the decomposed subtasks. GCP [27] and kSubS [7] use specialized subgoal predicting modules to search for plans. DHRL [19] uses efficient graph search by decoupling the time horizons of high-level and low-level policies.
While these methods have demonstrated success, they rely heavily on distance-based metrics, which are challenging to learn and sensitive to policy quality [10, 2]. In contrast, our method uses discrete reachability-based rewards that are easier to accurately estimate and provide clearer learning signals.
@@ 5 Discussion & Future Work
DHP architecture allows us to train a future-conditioned (goal) planning policy $\pi_{P}$ and a past-conditioned (memory) explorer policy $\pi_{E}$. The resulting model performs expertly on the standard 25-room long-horizon task, where the current SOTA approaches fail (Fig. 8(a)). The ablations show that the method generalizes beyond the training depths (Fig. 8(b)), the exploratory rewards significantly impact performance which are further enhanced using memory (Fig. 8(c)), and possible alternate reward schemes (Fig. 8(d)). Our method relies on imagination for training, which allows plan evaluation without the need to access the internal environmental state, but it may have inaccuracies (Fig. 9). However, we observed better results with discrete rewards than with the distance-based approach for our model (Fig. 8(d)). The architecture is flexible, and components can be used in isolation, e.g., the planning policy optimization can be combined with custom reachability measures. For future work, the agent can benefit from better and generic goal-state estimation mechanisms, preferably multimodal. The agent can also learn to generate goals for itself via a curriculum for targeted exploration, which can benefit exploration in complex environments with multiple bottleneck states that require a long-term commitment to a goal for efficient exploration. Given this, we show that the agent can be helpful at solving long-horizon tasks and believe that the ideas presented in the paper can be beneficial to the community. The code for the agent is available at: <URL redacted>.
@@ Appendix A Theoretical Proofs and Derivations
A.1 Policy Gradients for Trees
Given a goal-directed task as a pair of initial and final states $(s_{t},s_{g})$, a subgoal generation method predicts an intermediate subgoal $s_{0}$ that breaks the task into two simpler subtasks $(s_{t},s_{0})$ and $(s_{0},s_{g})$. The recursive application of the subgoal operator further breaks the task, leading to a tree of subtasks $\tau$ where each node $n_{i}$ is a task. Let the preorder traversal of the subtask tree $\tau$ of depth $D$ be written as $n_{0},n_{1},n_{2},...,n_{2^{D+1}-2}$. The root node $n_{0}$ is the given task and the other nodes are the recursively generated subtasks. Ideally, the leaf nodes should indicate the simplest reduction of the subtask that can be executed sequentially to complete the original task. The tree can be viewed as a trajectory where each node $n_{i}$ is a state, and taking action $\pi_{P}(z_{i}|n_{i})$ simultaneously places the agent in two states given by the child nodes $(n_{2i+1},n_{2i+2})$. Thus, the policy function can be written as $\pi_{P}(z_{i}|n_{i})$, the transition probabilities (can be deterministic also) as $p_{T}(n_{2i+1},n_{2i+2}|z_{i},n_{i})$, and the probability of the tree trajectory under the policy $\tau_{\pi_{P}}$ can be represented as:

### v3

@@ 2.4.1 Exploratory Reward
Since the planning policy $\pi_{p}$ uses the GCSR decoder for subgoal prediction, it is limited to the subgoals learned by the GCSR module. The GCSR training data must contain all possible state triplets in the environment for the planner to function well. To achieve this, we formulate our exploratory objective to encourage the Explorer to enact trajectories that are not well-modeled by the GCSR module. If Explorer traverses a trajectory that contains a state triplet $(s_{t},s_{t+q/2},s_{t+q})$, the modeling error is measured as the mean squared error between sub-goal $s_{t+q/2}$ and its prediction using the GCSR decoder. We compute the exploratory rewards for state $s_{t}$ based on the previous states as:
@@ 2.4.2 Memory Augmented Explorer
Since the exploratory rewards for the current step depend on the past states, the Explorer needs to know them to guide the agent accurately along rewarding trajectories. Fig. 5(c) shows the state dependencies for the exploratory rewards at different temporal resolutions $q$. Each temporal resolution $q$ requires $(s_{t-q},s_{t-q/2})$ for computing reward. For our case, the past states required are: $(s_{t-K},s_{t-2K}),(s_{t-2K},s_{t-4K}),$...] for each $q\in\{2K,4K,$...\} respectively. Removing the duplicates reduces the set of states required to: $\{s_{t},s_{t-K},s_{t-2K},$...\}.
@@
Figure 5: (a) Unconditional VAE that learns to predict states not conditioned on other states. (b) The explorer uses the memory and the VAE decoder to predict subgoals for the worker. (c) Illustration showing the states required as inputs for the Explorer for an example trajectory (top) being played out by an agent. It is a coarse trajectory that shows every $K$-th frame. The agent is at state $s_{t}$ and will receive rewards when it moves into the placeholder future state $s_{t+1}$ (dashed border). The rewards at $s_{t+1}$ will be computed using the GCSR for different temporal resolutions $q\in Q$ indicated on the right. The colored arrows indicate the state triplet required to compute the exploratory reward at $s_{t+1}$. Combining these state dependencies and removing redundancies yields the input requirements indicated below the dashed line. The inputs consist of the current state and the memory.
@@ 2.4.3 Policy Optimization
The Explorer is optimized as an SAC using the policy gradients from the REINFORCE objective. RSSM is used to imagine trajectories $\kappa$, followed by extracting every $K$-th frame of the rollout as abstract trajectories. For clarity, let the time steps of the abstract trajectory be indexed by $k$, where $k=t/K$. Then, the lambda returns $G^{\lambda}_{k}$ are computed for the abstract trajectories using a discount factor $\gamma_{L}$ (Eq. 11). Finally, the lambda returns are used to formulate the explorer actor and critic loss (Eq. 12,13).
@@ 3 Evaluation & Results
We extensively test our agent in the 25-room environment, a 2D maze task where the agent must navigate through connected rooms to reach the target position. Benchmarks from previous methods show average episode duration >150>150 steps, indicating a long-horizon task [pertsch2020long]. Observations are provided as initial and goal states ($64\times 64$ images) and a reward $0<R\leq 1$ upon reaching the goal position. Each episode lasts $400$ steps before terminating with a 0 reward. We follow the same evaluation protocols as the previous benchmarks ( [pertsch2020long]) and average across $100$ test runs. We also test our method in momentum-based environments, such as RoboYoga [lexa2021], which can be challenging due to the lack of momentum information in single images (the goal).
@@
We use a common hyperparameter setup for all tasks. The goal refresh rate is set to $K=8$, and the modeled temporal resolutions as $Q=\{2K,4K,8K,16K,32K\}$. The depth of the unrolled tree during training is $D=5$ and during inference is $D_{I}=8$ unless specified otherwise. For the first 33M steps, the explorer is used as the manager; then it shifts to the planning policy. The agent is trained every $16$ environmental steps. Please refer to section B for complete training details.
@@ 3.1 Results
Figure 7: Samples plan by our agent during training and inference. (a) Samples of full plans as sequences of subgoals from the start states (left) to the goal states (right). The subgoals are extracted from the goal states of all terminal nodes. The blanks occur when nodes terminate before maximum depth. (b) Sample subgoals generated during inference. The first and last images indicate the initial and goal states. Other images represent subgoals that break the path from the initial to the subgoal image on its right [Same order as 2(b)]. Blanks are when a reachable subgoal is found before the max depth.
We compare the performance in terms of the average success rate in reaching the goal state and the average path length against previous methods. Goal-Conditioned Behavioral Cloning (GC BC, [nair2017combining]) that learns goal-reaching behavior from example goal-reaching behavior. Visual foresight (VF, [ebert2018visual]) that optimizes rollouts from a forward prediction model via the cross-entropic method (CEM, [rubinstein2004cross, nagabandi2020deep]). Hierarchical planning using Goal-Conditioned Predictors (GCP,[pertsch2020long]) optimized using CEM to minimize the predicted distance cost. Goal-Conditioned Director (GC-Director). And LEXA, a SOTA sequential planning method that also uses an explorer and a planner but optimizes continuous rewards cosine and temporal. GC BC, VF, and GCP performances are taken from [pertsch2020long], which uses the same evaluation strategy. Table 1 shows that our model outperforms the previous work in terms of success rate and average episode lengths. Our method and LEXA (Temp) yield the shortest episode lengths. However, note that while LEXA explicitly optimizes for path lengths, our method uses an implicit objective.
Agent Success rate Average episode length Compute steps Time complexity
GC BC $7\%$ $402.48$ 11 11
GC-Director $9\%$ $378.89\pm 87.67$ 11 11
LEXA (Cos) $20\%$ $321.04\pm 153.29$ 11 11
VF $26\%$ $362.82$ MNMN NN
@@
Table 1: Average Performance of different approaches on the 25-room navigation task over $100$ evaluation runs. NN is the number of plan steps, and $M$ is the number of samples generated per plan.
Fig. 8(a) shows the score and episode length plots for some methods. Here, we plot an extra experiment, Director (Fixed Goal), where the goal remains fixed, and the agent only inputs the current state image. It can be seen that the agent shows signs of learning, the score falls but rises again around 88M steps to $\sim 70\%$. Comparing this to GC-Director (which completely fails) shows that the issue is not navigation or agent size, but the complexity of a goal-conditioned long-horizon task that requires planning.
@@
Figure 8: Experimental results (mean and stddev across 33 seeds) shown with episodic reward (top row) and average episode length (bottom row) against the environmental steps. The sharp rise in performance for DHP at step 3M indicates the switch from exploration to planning. (a) Baselines: Comparing our method DHP with state-of-the-art methods. (b) Bootstrapping: Compares shallow training depth models $(D=1,3)$ with the default training depth. (c) Explore data Comparison of planning policy performance trained on different data. (d) Comparison to Other reward schemes.
@@ 3.2 Ablations
We train two DHP agents with a maximum tree depth of $3$ and $1$ during training for planning. The $1$ depth agent is only allowed to break the given task once for learning. Fig. 8(b) shows the comparison of the agents, and it can be seen that all agents perform similarly. Note that LEXA with an equivalent horizon does not perform well. This demonstrates a key advantage of our bootstrapped return estimation: the value function learns to estimate returns for arbitrary depths, enabling the policy to plan beyond its training distribution. This is analogous to how TD learning in standard RL generalizes beyond observed trajectory lengths.
@@
We compare the planning policy trained against three data sources: (1) expert data collected using a suboptimal policy from [pertsch2020long](2) vanilla exploration using unconditional VAE reconstruction error, and (3) our memory-augmented GCSR-based exploration without memory (Sec. 2.4), and using an explorer without the memory. Fig. 8(c)shows our default memory-augmented explorer performs noticeably better, validating both the GCSR-based reward formulation and the memory mechanism. The superior performance over expert data is particularly noteworthy, as it demonstrates our task-agnostic exploration discovers more useful coverage than task-specific demonstrations.
Figure 9: Robo Yoga Environments
@@
We test our agent in the Deepmind Control Suite [tassa2018deepmind] based RoboYoga [lexa2021] environment. The task requires walker and quadruped agents to reach a goal body orientation specified by an image. The environment is challenging because it rewards the agent for maintaining the goal position, whereas our agent only plans to reach the goal state. Also, the goal states do not encode the momentum information. Fig. 9 shows the performance, of our agent at the tasks. Our agent can solve the walker task, but struggles at quadruped. We observed that our walker agent maintains the overall pose but constantly sways about the goal position. Fig. 13 shows an episode where the agent headstands constantly. Looking at the scaling limitations of our method, we implement a minimal version of DHP (Sec. F) and evaluate at the OGBench long-horizon large and giant mazes. The resulting agent outperforms the current best methods with an absolute improvement of $71.2\%$ in the giant humanoid maze (Tab. 4).
@@
We compare the default agent against a few variants: DHP (Neg Rew) That rewards 1-1 at non-terminal nodes and 0 at terminal (like an existence penalty), DHP (GAE) that estimates advantages as $\min$ of the GAE advantages for each child node, and DHP (Dist Sum) rewards negative estimated distances (between $n_{i,0}$ and $n_{i,1}$) at each node. Fig. 8(d) shows that the Neg Rew agent performs as well as the default method, while the others don’t perform as well.
@@ 4 Related Work
Hierarchical RL Agents: Hierarchical reinforcement learning (HRL) is a set of techniques that abstracts actions [botvinick2009hierarchically, wiering2012reinforcement, barto2003recent, sutton1999between, pateria2021hierarchical, nachum2018data]. Foundational works, such as the options framework [sutton1999between] and MAXQ decomposition [dietterich2000hierarchical], introduced temporal abstraction, enabling agents to reason at multiple time scales. Modern approaches learn hierarchical policies through mutual information (Causal InfoGAN [kurutach2018learning], DADS [sharma2020dynamics]), latent space planning (Director [hafner2022deep]), or trajectory encoding (OPAL [ajay2020opal]). These results demonstrate that hierarchical decomposition facilitates efficient credit assignment in planning.
Planning Algorithms: Planning methods aim to solve long-horizon tasks efficiently by exploring future states and selecting optimal actions [lavalle2006planning, choset2005principles]. Monte Carlo Tree Search (MCTS) [browne2012survey] expands a tree of possible future states by sampling actions and simulating outcomes. While effective in discrete action spaces, MCTS struggles with scalability in high-dimensional or continuous environments. Visual Foresight methods [ebert2018visual, finn2017deep, hafner2019learning] learned visual dynamics models to simulate future states, enabling planning in pixel-based environments. However, they require accurate world models and can be computationally expensive. Some use explicit graph search over the replay buffer data [eysenbach2019search]. Model Predictive Control (MPC) [nagabandi2018neural, nagabandi2020deep] is an online planner that samples future trajectories and optimizes actions over a finite horizon. These methods rely on sampling the future state and thus do not scale well with the horizon length. LEXA [kaelbling1993learning] is a policy-based linear planner that also uses an explorer for data collection.
To address the challenges of long-horizon tasks, some planning algorithms decompose complex problems into manageable subtasks by predicting intermediate subgoals [parascandolo2020divide, jurgenson2020sub, pertsch2020long]. MAXQ [dietterich2000hierarchical] decomposes the value function of the target Markov Decision Process (MDP) into an additive combination of smaller MDPs. Sub-Goal Trees [jurgenson2020sub] and CO-PILOT [ao2021co] learn a subgoal prediction policy optimized to minimize the total predicted distance measures of the decomposed subtasks. GCP [pertsch2020long] and kSubS [czechowski2021subgoal] use specialized subgoal predicting modules to search for plans. DHRL [lee2022dhrl] uses efficient graph search by decoupling the time horizons of high-level and low-level policies.
While these methods have demonstrated success, they rely heavily on distance-based metrics, which are challenging to learn and sensitive to policy quality [eysenbach2019search, ao2021co]. In contrast, our method utilizes discrete reachability-based rewards, which are easier to accurately estimate and provide clearer learning signals.
@@ 5 Discussion & Future Work
DHP architecture enables us to train a future-conditioned (goal) planning policy, $\pi_{P}$, and a past-conditioned (memory) exploration policy, $\pi_{E}$. The resulting model performs expertly on the standard 25-room long-horizon task than the current SOTA approaches (Fig. 8(a)). The ablations show that the method generalizes beyond the training depths (Fig. 8(b)), the exploratory rewards significantly impact performance, which are further enhanced using memory (Fig. 8(c)), and possible alternate reward schemes (Fig. 8(d)). Our method relies on imagination for training, which allows for plan evaluation without the need to access the internal environmental state; however, it may introduce inaccuracies (Fig. 9). However, we observed better results with discrete rewards than with the distance-based approach for our model (Fig. 8(d)). The architecture is flexible, and components can be used in isolation, e.g., the planning policy optimization can be combined with custom reachability measures. For future work, the agent can benefit from better and generic goal-state estimation mechanisms, preferably multimodal. The agent can also learn to generate goals for itself via a curriculum for targeted exploration, which can benefit exploration in complex environments with multiple bottleneck states that require a long-term commitment to a goal for efficient exploration. Given this, we demonstrate that the agent can be beneficial in solving long-range tasks and believe that the ideas presented in the paper can be valuable to the community independently. The code for the agent is available at: <URL redacted>.
@@ Appendix A Theoretical Proofs and Derivations
A.1 Notation Used
Table 2: Notation Summary
Symbol Description
States and Actions
$s_{t}$ State at time $t$ (RSSM representation)
$o_{t}$ Observation at time $t$ (raw image)
$s_{g}$ Goal state
$s_{wg}$ Worker goal (subgoal to be reached in $K$ steps)
$a_{t}$ Action at time $t$
$K$ Worker horizon (goal refresh rate, default: 8 steps)
Policies and Value Functions
$\pi_{W}(a_{t}|s_{t},s_{wg})$ Worker policy (low-level action selection)
$\pi_{E}(z|s_{t},\text{mem}_{t})$ Explorer policy (goal generation for exploration)
$\pi_{P}(z|s_{t},s_{g})$ Planning policy (subgoal generation for task)
$v_{W}(s_{t},s_{wg})$ Worker value function
$v_{E}(s_{t},\text{mem}_{t})$ Explorer value function
$v_{P}(n_{i})$ Planning policy value function (for tree node $n_{i}$)
Goal-Conditioned State Recall (GCSR)
$\text{Enc}_{G}(z|s_{t},s_{t+q/2},s_{t+q})$ GCSR encoder (outputs latent distribution)
$\text{Dec}_{G}(s_{t},s_{t+q},z)$ GCSR decoder (predicts midway state)
$z$ Latent variable in GCSR (4×4 mixture of categoricals)
$Q$ Set of temporal resolutions, $\{2K,4K,8K,16K,32K\}$
$q\in Q$ Temporal resolution (distance between states in triplet)
Unconditional VAE (Explorer)
$\text{Enc}_{U}(z|s_{t})$ Unconditional encoder (for explorer)
$\text{Dec}_{U}(z)$ Unconditional decoder (for explorer)
$\text{mem}_{t}$ Memory buffer (past states: $\{s_{t-K},s_{t-2K},\ldots\}$)
Tree Planning
$\tau$ Tree trajectory (complete subtask tree)
$n_{i}$ Node $i$ in tree trajectory (represents a subtask)
$n_{i,0}$ Initial state of subtask at node $i$
$n_{i,1}$ Goal state of subtask at node $i$
$T_{i}\in\{0,1\}$ Terminal indicator (1 if node $i$ is reachable by worker)
$R_{i}\in\{0,1\}$ Reward at node $i$ (1 if terminal, 0 otherwise)
$G_{i}^{\lambda}$ Lambda return for node $i$ (Eq. 6)
$D$ Maximum tree depth during training (default: 5)
$D_{I}$ Maximum tree depth during inference (default: 8)
$\Delta_{R}$ Reachability threshold for cosine similarity (default: 0.7)
Discount Factors and Hyperparameters
$\gamma$ Discount factor for tree returns (default: 0.95)
$\gamma_{L}$ Discount factor for linear returns—worker & explorer (default: 0.99)
$\lambda$ Lambda parameter for TD($\lambda$) returns (default: 0.95)
$\beta$ KL loss weight for VAE training (default: 1.0)
$\eta$ Entropy coefficient for policy training (default: 0.5)
Offline Variant (Section F)
$\phi(s_{t},s_{g})$ Goal representation function (length-normalized vector)
$V^{h}(s_{t},s_{g})$ High-level (manager) value function
$V^{l}(s_{t},z)$ Low-level (worker) value function
$\pi_{M}(z|s_{t},s_{g})$ Manager policy (predicts goal representations)
$\mathcal{B}$ Goal buffer (stores diverse landmark states)
$\mathcal{B}(s_{t},z)\rightarrow s_{g}$ Buffer retrieval (nearest state to representation $z$)
$\tilde{V}^{l}$ Normalized value: $(V^{l}-\mu_{V^{l}})/\sigma_{V^{l}}$
$\theta$ Reachability threshold for normalized value (default: 2.0-2.0)
$\tau$ Expectile parameter for IQL-style training (default: 0.7)
$\gamma^{h}$ High-level discount factor (default: 0.9)
A.2 Policy Gradients for Trees
Given a goal-directed task as a pair of initial and final states $(s_{t},s_{g})$, a subgoal generation method predicts an intermediate subgoal $s_{0}$ that breaks the task into two simpler subtasks $(s_{t},s_{0})$ and $(s_{0},s_{g})$. The recursive application of the subgoal operator further breaks down the task, leading to a tree of subtasks $\tau$, where each node $n_{i}$ represents a task. Let the preorder traversal of the subtask tree $\tau$ of depth $D$ be written as $n_{0},n_{1},n_{2},...,n_{2^{D+1}-2}$. The root node $n_{0}$ is the given task, and the other nodes are the recursively generated subtasks. Ideally, the leaf nodes should indicate the simplest reduction of the subtask that can be executed sequentially to complete the original task. The tree can be viewed as a trajectory, where each node $n_{i}$ represents a state, and taking action $\pi_{P}(z_{i}|n_{i})$ simultaneously places the agent in two states, given by the child nodes $(n_{2i+1},n_{2i+2})$. Thus, the policy function can be written as $\pi_{P}(z_{i}|n_{i})$, the transition probabilities (can be deterministic also) as $p_{T}(n_{2i+1},n_{2i+2}|z_{i},n_{i})$, and the probability of the tree trajectory under the policy $\tau_{\pi_{P}}$ can be represented as: